OpenStudy (anonymous):

Multiply: 4x^y/12x x 24x^2y^2/6xy

OpenStudy (anonymous):

\[\frac{4x^y}{12x}\times\frac{24x^2y^2}{6xy}\]?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Oh wait no. Hold on.

OpenStudy (anonymous):

\[\frac{4x ^{2}y }{12x } \times \frac{24x^{2}y^{2} }{6xy}\]

OpenStudy (anonymous):

@Peter14 ^^

OpenStudy (anonymous):

ok, that makes more sense. Just multiply straight across and simplify

OpenStudy (anonymous):

So \[96x ^{4}y ^{2}\] first line?

OpenStudy (anonymous):

@Peter14

OpenStudy (anonymous):

96 x^4 y^3

OpenStudy (anonymous):

Oh oops. See my mistake.

OpenStudy (anonymous):

\[72x^{2}y\] second line? @Peter14

OpenStudy (anonymous):

looks right

OpenStudy (anonymous):

\[\frac{ 96x^{4}y^{3} }{ 72x^{2}y }\]

OpenStudy (anonymous):

Now do I simplify?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay so it goes to \[\frac{ 4 }{ 3 }\] and then..?

OpenStudy (anonymous):

you still have x^2 y^2

OpenStudy (anonymous):

How do I reduce those?

OpenStudy (anonymous):

Can you finish it for me? I still have other questions to ask and it's late. Please & thank you.

OpenStudy (anonymous):

you can represent 96 x^4 y^3/72 x^2 y as 

OpenStudy (anonymous):



OpenStudy (anonymous):

cancel xs with xs and ys with ys (this works because anything/itself is zero) and you simplify 96/72 to 4/3

OpenStudy (anonymous):

That makes it much clearer thanks! So \[\frac{ 96x ^{2}y }{ 72 }\]

OpenStudy (anonymous):

4x^2y / 3

OpenStudy (anonymous):

Appreciate it!

OpenStudy (anonymous):

yes, you're right

hartnn (hartnn):

wouldn't it be y^2 in numerator ?

hartnn (hartnn):

4x^2y^2 / 3

OpenStudy (anonymous):

oh, yes, thank you hartnn

OpenStudy (anonymous):

Thanks!

hartnn (hartnn):

welcome ^_^